Grade 11Physical Chemistry

How to calculate the total number of protons,electrons,& neutrons present in 14g of C-14 is

Lakshmi
8 Years agoGrade 11

1 Answer

Purva
8 Years ago
number of moles=1 (given weight / molecular weight)
molecules = moles*NA = 1*NA (NA=6.023*10^23 )
  One atom of carbon-14 (atomic number=6) contains 6 protons, 6 electrons and 14-6=8 neutrons so in 14g,ie,1  mole:
number of protons =6*NA
number of electrons= 6*NA
number of neutrons=(14-6)*NA = 8*NA
Please approve if you like my answer....